OpenStudy (kpop_fan):

A box contains 100 colored chips; some are yellow and some are green. John chooses a chip at random, records the color, and places it back in the bag. John has recorded 57 yellow chips and 19 green chips. Using these results, what is the predicted number of green chips in the box? 25 43 76 81

OpenStudy (anonymous):

Step 1: add 57 and 19

OpenStudy (anonymous):

25 ? i added both 57 and 19 and then subtracted from 100

OpenStudy (anonymous):

Step 2: divide the number of green chips by 76

OpenStudy (anonymous):

Step 2: divide the number of green chips by 76

OpenStudy (kpop_fan):

57

OpenStudy (anonymous):

you got this ... great job, keep it up

OpenStudy (mrnood):

@okay101 your approach is not correct - I suggest oyu leave it to @MARC_

OpenStudy (kpop_fan):

sorry I meant 25

OpenStudy (anonymous):

Step 1: add 57 and 19\[57+19=76\]Step 2:divide the number of green chips by 76\[\frac{ 19 }{ 76 }= ?\]

OpenStudy (anonymous):

what value will u get is u divide 19 by 76?

OpenStudy (anonymous):

if*

OpenStudy (kpop_fan):

19/76= 0.25 so its 25?

OpenStudy (anonymous):

yeah i am sorry i thought that i was doing something different, i apologize @MrNood @MARC_

OpenStudy (anonymous):

correct! The answer is 25...

OpenStudy (kpop_fan):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i was right @MrNood ... haha

OpenStudy (mrnood):

you were right - I didn't say you weren't - but your METHOD was absolutely wrong. It is important to post HELP not answers - as just '25' does not teach the poster anything at all

OpenStudy (anonymous):

well , people do things different ways @MrNood

OpenStudy (mrnood):

I appreciate that there are sometimes more than 1 way to tackle a maths problem. However - in this case YOUR way is not a valid method , and your arrival at the correct answer was just luck